The force on a wire carrying 8.45 A is a maximum of 1.78 N when placed between the pole faces of a magnet.

If the pole faces are 51.5 cm in diameter, what is the approximate strength of the magnetic field?
Answer #1

We know that magnetic force on the current carrying wire

F=1(L x B) = ILB sin(0)

Here theta =900 , I=8.45A , F=1.78 N ,

L=51.5cm=0.515 m ( since 0.515 m wire in the magnetic field which is diameter of pole)

put these value in equation 1

1.78=8.45×0.515×B×sin(90)

B=1.78/(0.515×8.45)

B=0.409 T. Answer